# EtCon: Edit-then-Consolidate for Reliable Knowledge Editing

### v1

@@ Abstract
Knowledge editing aims to update specific facts in large language models (LLMs) without full retraining. Prior efforts sought to tune the knowledge layers of LLMs, proving effective for making selective edits. However, a significant gap exists between their performance in controlled, teacher-forcing evaluations and their real-world effectiveness in lifelong learning scenarios, which greatly limits their practical applicability. This work’s empirical analysis reveals two recurring issues associated with this gap: (1) Most traditional methods lead the edited model to overfit to the new fact, thereby degrading pre-trained capabilities; (2) There is a critical absence of a knowledge consolidation stage, leaving new facts insufficiently integrated into LLMs’ inference-time behavior under autoregressive generation, thereby leading to a mismatch between parametric knowledge and actual generation behavior. To this end, we propose Edit-then-Consolidate, a novel knowledge editing paradigm that aims to bridge the gap between theoretical knowledge editing methods and their real-world applicability. Specifically, (1) our framework mitigates overfitting via Targeted Proximal Supervised Fine-Tuning (TPSFT) that localizes the edit via a trust-region objective to limit policy drift; (2) Then, a consolidation stage using Group Relative Policy Optimization (GRPO) aligns the edited knowledge with CoT-based inference policy by optimizing trajectory-level behavior under comprehensive reward signals. Extensive experiments demonstrate our framework consistently improves editing reliability and generalization under real-world evaluations, while better preserving locality and pre-trained capabilities.
@@ 1 Introduction
Large language models (LLMs) have demonstrated unprecedented capabilities across numerous tasks Guo et al. (2025); Wang et al. (2025b), serving as foundational reasoning engines for information retrieval Yang et al. (2025a), task automation agents He et al. (2025); Liu et al. (2025b); Wang et al. (2025a), and scientific research Rosen et al. (2025); Shmatko et al. (2025). However, as the external world continuously evolves, the static nature of LLMs’ pre-trained knowledge renders specific versions rapidly obsolete Zheng et al. (2025a). While retraining a large-scale LLM with updated knowledge could address this limitation, it requires substantial computational resources and pre-training data to maintain both knowledge update efficacy and general capabilities, making frequent knowledge updates impractical Mitchell et al. (2022). Knowledge editing methods Zhang et al. (2025); Scialanga et al. (2025); Li et al. (2025b); Rozner et al. (2024) have thus garnered significant attention as techniques that achieve targeted knowledge updates through localized parameter modifications while avoiding extensive resource consumption.
Knowledge editing methods can be categorized into three main paradigms: (1) Parametric in-place editing methods Meng et al. (2022a); Dai et al. (2025); Han et al. (2024); Fu et al. (2025), which directly compute weight updates and apply them to the LLM’s weight matrices, encompassing approaches such as locate-then-edit, parameter-efficient fine-tuning, and model merging—all of which preserve model architecture without requiring additional modules; (2) Meta-learning-based methods Hartvigsen et al. (2023); Li et al. (2025b); Tan et al. (2023) that train auxiliary hypernetworks to predict weight updates for specific parameters to achieve knowledge editing objectives; (3) Memory-based methods Wang et al. (2024d; a) that store new knowledge in external modules and train LLMs to activate these modules during inference involving updated knowledge. While these methods demonstrate promise in constrained evaluation scenarios such as single editing and teacher-forcing evaluations, a significant performance gap emerges in more realistic auto-regressive evaluation and lifelong editing Jiang et al. (2024); Tan et al. (2023); Chen et al. (2024), leading recent research Gu et al. (2024a); Huang et al. (2024) to question the reliability and practical utility of existing knowledge editing methods.
In this work, we conduct a comprehensive investigation into the root causes of this performance gap, focusing on Parametric In-Place Editing methods due to their high potential for practical application in lifelong learning scenarios. Through comprehensive empirical analysis, we identify two critical issues at the root of this gap. First, most of traditional methods cause edited models to overfit to newly introduced facts. This overfitting leads to excessive specialization of model parameters to editing examples, thereby degrading pre-trained general capabilities including robust reasoning, linguistic fluency, and robustness. Second, and more critically, a fundamental absence of a dedicated knowledge consolidation phase is observed. This omission results in new information being superficially encoded at the parametric level, failing to establish deep integration with the LLM’s inference-time behavior under autoregressive generation. This discordance manifests as a critical decoupling between knowledge representation and inferential activation. While the model successfully incorporates updated knowledge parametrically, it consistently fails to retrieve, activate, or apply this knowledge during autoregressive generation.
Figure 1 illustrates the core challenge we address. When existing methods edit a fact (e.g., updating Michael Jordan’s nationality), the model may parametrically encode the new information but fail to consistently apply it during reasoning. This manifests as contradictory outputs where the model simultaneously acknowledges both old and new facts, revealing a fundamental misalignment between parametric knowledge and actual generation behavior. Our Edit-then-Consolidate framework resolves this by introducing a crucial consolidation stage that aligns the edited knowledge with the model’s inference-time policy. To address the limitation illustrated above, we propose Edit-then-Consolidate (EtCon), a two-stage knowledge-editing paradigm. In the first stage, we employ Targeted Proximal Supervised Fine-Tuning (TPSFT)—a refined variant of PSFT Zhu et al. (2025) that selectively updates only the FFN layers identified as knowledge repositories. This targeted approach, combined with trust-region constraints, ensures localized edits that preserve the model’s broader capabilities. In the second stage, we introduce a critical consolidation phase using Group Relative Policy Optimization (GRPO) to align the parametric knowledge with the model’s CoT-based inference policy through trajectory-level optimization under comprehensive reward signals.
@@
We conduct extensive experiments on three datasets with Llama-3-8B-Instruct and Qwen2.5-7B-Instruct. Under auto-regressive generation with natural stopping and an LLM-as-a-judge protocol using GPT-4.1 (OpenAI (2025)), Edit-then-Consolidate improves editing reliability and generalization by 35%-50% over strong baselines. It also significantly enhances locality while preserving critical pre-trained capabilities. Our contributions can be summarized as follows: (1) We identify that the absence of a knowledge-consolidation stage creates a critical knowledge-behavior misalignment, serving as the key bottleneck to the real-world applicability of knowledge-editing methods. (2) We propose Edit-then-Consolidate (EtCon): TPSFT for localized parametric edits, followed by GRPO for trajectory-level consolidation that aligns parametric knowledge with actual generation behavior. (3) Extensive experiments demonstrate that EtCon improves editing reliability and generalization by 40%–50%, strengthens locality, and preserves pre-trained capabilities under realistic evaluation settings.
@@ 2.1 Overview of Knowledge editing methods
From the perspective of model architecture, the three paradigms introduced above can also be coarsely grouped into two families. Parametric in-place editing methods preserve the vanilla LLM architecture. The locate-then-edit paradigm Meng et al. (2022a); Dai et al. (2025); Li et al. (2024); Zhong et al. (2025); Zhang et al. (2024c) identifies knowledge locations within LLMs and modifies targeted parameters through gradient-based or analytical solutions. PEFT methods Zhu et al. (2020); Han et al. (2024); Wang et al. (2024c); Gupta et al. (2025) directly update model parameters via regularized gradient descent to achieve knowledge updates while constraining side effects Liu et al. (2025a). These approaches seamlessly integrate with existing deployment infrastructure without additional inference latency. External-assisted editing methods rely on auxiliary modules for knowledge modification. Meta-learning approaches Tan et al. (2023); Hartvigsen et al. (2023); Li et al. (2025b) train hypernetworks to generate parameter updates, while memory-based methods Hartvigsen et al. (2023); Zhang et al. (2024b); Chen et al. (2024) encode knowledge in external modules that the LLM retrieves during inference. Despite their superior performance in balancing reliability and locality, external methods introduce deployment complexity. Given these trade-offs, our work advances parametric in-place editing for lifelong knowledge editing scenarios.
@@ 2.2 Evaluation of Knowledge editing methods
Existing research Fang et al. (2024); Qi et al. (2025); Scialanga et al. (2025) predominantly evaluates the effectiveness of knowledge editing methods using a standard set of metrics. Reliability assesses the success rate of editing by calculating the percentage where . Generalization evaluates the model’s ability to generalize to new knowledge post-editing, measured by the percentage where when presented with rephrased queries pertaining to the new knowledge. Locality measures the extent to which editing a specific fact preserves the model’s responses to questions related to neighboring, unedited facts. Conventionally, the evaluation input typically consists of simple queries with identical prompt formats, without additional contextual information. For the output, edited models’ responses are often truncated to a target answer length or constrained by examples to match a specific target answer format. In generation, teacher forcing is frequently employed, feeding ground truth tokens as input during the decoding process. Recent studies, however, have highlighted the fragility of such evaluation paradigms. Consequently, this paper adopts a realistic evaluation approach for knowledge editing methods. Details of the real-world evaluation framework are in appendix A.7
3 The Missing Consolidation Stage in Knowledge Editing
Recent studies have revealed a critical performance gap in knowledge editing methods: while achieving high success rates under controlled teacher-forcing evaluation, these methods exhibit catastrophic failures in realistic auto-regressive settings. This stark discrepancy undermines their practical utility and raises fundamental questions about the effectiveness of current approaches. Through systematic investigation, we identify that this failure stems not from the editing mechanism itself, but from a fundamental architectural omission—the absence of a knowledge consolidation stage. We hypothesize that successful knowledge updating requires a two-stage process: (1) an initial parametric edit that injects new information into LLMs’ weights, followed by (2) a consolidation phase that integrates this knowledge into the LLMs’ inference-time behavior under autoregressive generation. Without consolidation, edited knowledge remains superficially encoded at the parametric level, failing to propagate to the model’s actual generation behavior.
| Method | Reli. | Gener. | Local. | |
Llama-3-8b-Instruct |
Pre-Edit | 2.8 | 2.4 | 38.6 |
| Pre-Edit(+GRPO) | 5.2 | 4.7 | 38.4 | |
| FT-M | 16.6 | 15.5 | 29.3 | |
| FT-M(+GRPO) | 62.9 | 52.7 | 24.9 | |
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | |
| ALPHAEDIT(+GRPO) | 50.4 | 38.7 | 5.4 |
To validate this hypothesis, we conduct controlled experiments augmenting existing editing methods with a consolidation mechanism. Table 1 presents compelling evidence: introducing Group Relative Policy Optimization (GRPO) as a post-editing consolidation step dramatically transforms performance. For FT-M, reliability surges from 16.6% to 62.9% on Llama3-8b-Instruct, while ALPHAEDIT improves from 18.7% to 50.4%. Crucially, these gains extend to generalization metrics, indicating that consolidation enables genuine knowledge integration rather than superficial memorization. The reward trajectories in Fig. 2 further illuminate the consolidation dynamics. The monotonic increase demonstrates stable knowledge integration, where the model progressively aligns its reasoning behavior with the edited knowledge. Notably, applying GRPO directly to unedited models yields minimal improvements (Pre-Edit: 2.8% → 5.2%), confirming that consolidation requires prior parametric editing as a foundation. These findings establish a critical insight: the limitations of current knowledge editing methods arise from treating editing as a single-stage process. The Edit-then-Consolidate paradigm we propose addresses this fundamental gap, recognizing that parametric updates and behavioral alignment are complementary but distinct requirements for successful knowledge editing.
4 The Edit-then-Consolidate Framework
Building on the observational evidence in the preceding section, we posit that the limitations of current LLM knowledge-editing methods arise primarily from the lack of a principled consolidation stage that integrates edited knowledge with the model’s reasoning behavior; moreover, repeated overfitting edits can erode general abilities. To address this, we introduce Edit-then-Consolidate paradigm: Stage I employs Targeted Proximal Supervised Fine-Tuning (TPSFT) to perform localized knowledge editing under trust-region–style constraints, thereby limiting spillover while preserving pre-trained abilities; Stage II applies Group Relative Policy Optimization (GRPO) with a task-appropriate comprehensive reward to consolidate at the trajectory level under real-world evaluation signals. The remainder of this section presents the design rationale and the interaction between these two stages.
@@ 4.1 Knowledge Editing via Targeted Proximal Fine-Tuning
In this section, we introduce Targeted Proximal Supervised Fine-Tuning (TPSFT) as a refined knowledge-editing method that addresses the trilemma of reliability, locality, and generality. This approach differs from raw PSFT that update the whole LLMs, by selectively update only the FFNs of LLMs. This targeted update strategy effectively injects new knowledge while minimizing disruption to the model’s overall architecture and pre-trained capabilities. We consider a knowledge editing dataset that contains editing instances , where each context contains a question about the new fact, and is the corresponding ground-truth answer. At the start of the editing process, we have a vanilla LLM parameterized by . We partition the model’s parameters into two disjoint sets: the target FFN parameters to be edited, , and the remaining frozen parameters, , such that . The objective of TPSFT is to learn a new set of FFN parameters, , yielding an updated model where . This model must accurately generate the target answer for a given context , while minimizing disruptions to its performance on unrelated inputs. A critical innovation in our TPSFT is the use of CoT-augmented training labels. For each editing instance , we: (1) prompt the vanilla LLM to generate a CoT reasoning path for using an instruction template (see Appendix), and (2) replace the generated answer with the target new fact , yielding the training label . This design enables learning smoothed distributions over reasoning paths rather than sharp one-hot targets. More importantly, it preserves the model’s natural reasoning patterns—the model learns to reach new answers through its inherent reasoning style rather than abandoning pre-trained capabilities. This significantly reduces disruption while ensuring accurate knowledge updates.
To achieve this, we update the targeted FFN parameters by minimizing the following Targeted Proximal Supervised Fine-Tuning (TPSFT) loss over the editing dataset , while the rest of the model parameters remain frozen:
@@
Here, is a hyperparameter that defines the clipping radius, which controls the size of the trust region. The probability ratio is the core of this objective and is defined as:
@@
where is the probability of generating the ground-truth answer given the context from the model with updated FFN parameters, and is the corresponding probability from the reference policy. At the start of the editing process, this reference policy is the initial vanilla LLM. For each subsequent edit instance in the sequential editing process, it is then updated to be the state of the model resulting from the immediately preceding edit.
This objective function creates a trust-region constraint that is critical for balanced knowledge editing. The term aims to increase the likelihood of the correct answer, which is analogous to the objective in standard supervised fine-tuning. However, the ‘clip‘ function prevents this ratio from deviating too far from 1. When the updated model becomes significantly more confident about the target answer than the original model (i.e., when ), the gradient signal is effectively nullified for that instance. This mechanism acts as a powerful regularization, discouraging overly aggressive updates that could lead to overfitting on the new fact and, consequently, the disruption of pre-trained capabilities.
By integrating targeted parameter updates with a constrained optimization objective, TPSFT directly addresses the editing trilemma. Locality is achieved by physically confining the updates to the FFN layers, which are hypothesized to be the primary repositories of factual knowledge. Reliability is enforced by the supervised objective that maximizes the probability of the new fact. Finally, stability is preserved by the PSFT clipping mechanism, which prevents drastic policy shifts and ensures that the model’s behavior remains stable and consistent across a wide range of inputs beyond the specific edit.
4.2 Knowledge Consolidation via Group Relative Policy Optimization
After the TPSFT stage, the edited model has incorporated new facts at the parametric level. However, these parametric changes do not automatically propagate to the model’s reasoning capabilities. To bridge this gap, we introduce a consolidation step using Group Relative Policy Optimization (GRPO) that aligns the model’s inference-time behavior with the injected knowledge.
We formulate the consolidation as a reinforcement learning problem. Given a reasoning dataset containing queries that require reasoning over the edited facts, we optimize the edited model to generate trajectories that demonstrate both factual accuracy and reasoning consistency. The objective maximizes expected reward while constraining deviation from the post-TPSFT model:
where serves as the reference policy (the model after TPSFT), and controls the strength of regularization.
We optimize this objective using the GRPO algorithm with the following surrogate loss:
where is the importance ratio, and is the group-relative advantage computed from a batch of sampled trajectories.

### v2

# EtCon: Edit then Consolidate for Reliable Knowledge Editing

@@ Abstract
Knowledge editing aims to update specific facts in large language models (LLMs) without full retraining. Prior efforts sought to tune the knowledge layers of LLMs, achieving improved performance in controlled, teacher-forced evaluations. However, they still encounter challenges in real-world autoregressive generation scenarios, which greatly limit their practical applicability. Our empirical analysis reveals two issues: (1) Most methods degrade pre-trained capabilities after injecting new knowledge; (2) They may exhibit a discrepancy between stored parametric knowledge and inference-time autoregressive generation behavior. To this end, we propose EtCon, an edit-then-consolidate paradigm that couples targeted edits with post-edit consolidation. Specifically, our framework comprises two stages: (1) Targeted Proximal Supervised Fine-Tuning (TPSFT) performs a constrained targeted edit to update parametric knowledge while controlling policy drift. (2) Group Relative Policy Optimization (GRPO) consolidates the edit by aligning autoregressive trajectories with the intended fact. Extensive experiments demonstrate that our EtCon improves editing reliability and real-world generalization, while better preserving pre-trained capabilities.
[…]
@@ 1 Introduction
Large language models (LLMs) have demonstrated unprecedented capabilities across numerous tasks (Wang et al., 2025; Yang et al., 2025a; He et al., 2025; Liu et al., 2025b; Rosen et al., 2025; Shmatko et al., 2025). However, as the world continuously evolves, static pre-trained knowledge quickly becomes outdated (Zheng et al., 2025a). While continual pre-training can refresh knowledge, it is computationally expensive and impractical to perform frequently at scale (Mitchell et al., 2022), motivating knowledge editing methods that perform targeted updates through localized parameter modifications (Zhang et al., 2025; Scialanga et al., 2025; Li et al., 2025; Rozner et al., 2024). Existing knowledge editing methods roughly fall into two families: In-Place Parametric Edits (Meng et al., 2022a; Dai et al., 2025; Han et al., 2024; Fu et al., 2025; Jiang et al., 2025) that directly update model weights under the vanilla architecture, and External-Assisted Edits (Hartvigsen et al., 2023; Li et al., 2025; Tan et al., 2023; Wang et al., 2024d, a) that introduce auxiliary modules (e.g., hypernetworks or memory) to support updates at inference time. Despite encouraging results under single-edit and teacher-forced evaluations, their reliability in real-world autoregressive generation remains limited (Jiang et al., 2024; Tan et al., 2023; Chen et al., 2024; Gu et al., 2024a; Huang et al., 2024).
In this work, our empirical analysis reveals two recurring issues. (1) Pre-trained capability degradation. Many existing methods perform edits by maximizing the likelihood of the new target on a small set of editing samples. Without explicit constraints or regularization, this few-shot objective can induce over-specialization to the editing examples, degrading pre-trained capabilities. In lifelong editing, as illustrated in Figure 1 (left), such drift may accumulate across lifelong edits and may even lead to model collapse (Gupta et al., 2024, 2025b). (2) Parametric knowledge–behavior mismatch in autoregressive generation. More critically, Figure 1 (right) shows that even when the edited fact is stored parametrically, it may not be consistently expressed along autoregressive trajectories. A key reason is the train-test mismatch between teacher forcing and autoregressive inference: edits are optimized on ground-truth prefixes, whereas real-world generation conditions on self-generated prefixes, inducing a prefix distribution shift that makes edited behavior fragile and prone to reversion.
@@
To address these limitations, we propose EtCon, a new edit-then-consolidate paradigm (Fig. 4) that couples constrained targeted edits with post-edit consolidation to improve both editing reliability and pre-trained capability preservation. Our training involves a two-stage pipeline: Stage 1 (Edit). Targeted Proximal Supervised Fine-Tuning (TPSFT), a refined variant of PSFT (Zhu et al., 2025), performs a constrained targeted edit by updating only knowledge-relevant FFN layers, with trust-region constraints and self-generated CoT-based targets. This design localizes parameter updates and mitigates over-specialization, thereby preserving pre-trained capabilities. Stage 2 (Consolidate). GRPO consolidates the edit via trajectory-level optimization under reward signals. This stage aligns the parametric update with inference-time behavior, improving the robustness of the edited knowledge along autoregressive trajectories. We conduct extensive experiments on three datasets with Llama-3-8B-Instruct and Qwen2.5-7B-Instruct. Under auto-regressive generation with natural stopping and an LLM-as-a-judge protocol (OpenAI, 2025), EtCon improves editing reliability and generalization by 35%-50% over strong baselines. It also significantly enhances locality while preserving critical pre-trained capabilities. Our contributions are as follows:
(1) We empirically show that incorporating a post-edit consolidation stage improves knowledge editing reliability under realistic scenarios.
(2) We propose EtCon, a novel knowledge editing paradigm: TPSFT for localized parametric edits, followed by GRPO for trajectory-level consolidation that aligns parametric knowledge with actual generation behavior.
(3) Extensive experiments demonstrate that EtCon improves editing reliability and generalization, while preserving pre-trained capabilities under realistic evaluation settings.
| Method | Reli. | Gener. | Local. | |
| Llama-3-8B-Instruct | Pre-Edit | 2.8 | 2.4 | 38.6 |
| Pre-Edit(+GRPO) | 5.2 (+2.4) | 4.7 (+2.3) | 38.4 (-0.2) | |
| FT-M | 16.6 | 15.5 | 29.3 | |
| FT-M(+GRPO) | 62.9 (+46.3) | 52.7 (+37.2) | 24.9 (-4.4) | |
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | |
| ALPHAEDIT(+GRPO) | 50.4 (+31.7) | 38.7 (+24.7) | 5.4 (-0.9) |
@@ 2.1 Overview of Knowledge editing methods
From the perspective of model architecture, the three paradigms introduced above can also be coarsely grouped into two families. In-Place Parametric Edits methods preserve the vanilla LLM architecture. The locate-then-edit paradigm (Meng et al., 2022a; Dai et al., 2025; Li et al., 2024; Zhong et al., 2025; Zhang et al., 2024c) identifies knowledge locations within LLMs and modifies targeted parameters through gradient-based or analytical solutions. PEFT methods (Zhu et al., 2020; Han et al., 2024; Wang et al., 2024c; Gupta et al., 2025a) directly update model parameters via regularized gradient descent to achieve knowledge updates while constraining side effects (Liu et al., 2025a). These approaches seamlessly integrate with existing deployment infrastructure without additional inference latency. External-Assisted Edits methods rely on auxiliary modules for knowledge modification. Meta-learning approaches (Tan et al., 2023; Hartvigsen et al., 2023; Li et al., 2025) train hypernetworks to generate parameter updates, while memory-based methods (Hartvigsen et al., 2023; Zhang et al., 2024b; Chen et al., 2024) encode knowledge in external modules that the LLM retrieves during inference. Despite their superior performance in balancing reliability and locality, external methods introduce deployment complexity. Given these trade-offs, our work advances In-Place Parametric Edits with a focus on more realistic lifelong knowledge editing scenarios.
@@ 2.2 Evaluation of Knowledge editing methods
Existing work (Fang et al., 2024; Qi et al., 2025; Scialanga et al., 2025) typically evaluates knowledge editing with three standard metrics. Reliability measures edit success as the fraction of cases where . Generalization applies the same criterion to rephrased queries about the edited knowledge. Locality assesses whether the edit preserves responses to related, unedited facts. In these protocols, inputs are often templated queries with minimal context, outputs are constrained to a fixed answer format (e.g., via truncation or in-context exemplars), and decoding commonly uses teacher forcing with ground-truth prefixes. Recent studies, however, have highlighted the fragility of such evaluation paradigms. Consequently, this paper adopts a realistic evaluation approach for knowledge editing methods. The full real-world evaluation framework is described in Appendix A.6.
3 Preliminary
3.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) (Guo et al., 2025) is a trajectory-level reinforcement learning algorithm that optimizes language model policies through relative reward comparisons within sampled groups. Given a policy and a dataset , GRPO maximizes expected reward while constraining deviation from a reference policy :
where is a reward function and controls regularization strength. The optimization employs a clipped surrogate objective:
where is the importance ratio, and is the group-relative advantage computed from trajectories sampled for the same query . This design eliminates the need for a separate critic network while enabling stable policy updates. Notably, GRPO’s trajectory-level optimization makes it naturally suited for aligning knowledge that is parametrically encoded but not reliably applied during autoregressive generation.
3.2 Motivation for Knowledge Consolidation
Our key insight is that successful knowledge editing requires a two-stage process rather than parametric knowledge editing alone. Specifically, recent studies reveal a critical gap: while existing methods achieve high success under teacher-forcing evaluation, they can fail in realistic autoregressive generation. Through systematic investigation, we posit that an explicit knowledge consolidation stage is crucial for mitigating this gap. Without a knowledge consolidation stage, edited knowledge may remains superficially encoded at the parametric level, failing to propagate to the model’s generation behavior. To validate this hypothesis, we augment existing editing methods with GRPO as a post-editing consolidation step. As shown in Table 3 and Figure 3, introducing consolidation dramatically improves both reliability and generalization across multiple baselines. Furthermore, applying GRPO directly to unedited models yields minimal gains, confirming that consolidation requires prior parametric editing as a foundation. These findings establish that parametric knowledge updates and behavioral alignment are complementary but distinct requirements for successful knowledge editing.
4 EtCon
In this section, we present the EtCon framework (as shown in Figure 4). Stage I employs Targeted Proximal Supervised Fine-Tuning (TPSFT) to perform localized knowledge editing with PPO-style ratio clipping that constrains policy divergence, thereby limiting spillover while preserving pre-trained abilities (§4.1); Stage II applies GRPO with a comprehensive reward to consolidate edited knowledge at the trajectory level under autoregressive generation (§4.2).
@@ 4.1 Knowledge Editing via Targeted Proximal Fine-Tuning
In this section, we introduce Targeted Proximal Supervised Fine-Tuning (TPSFT) as a refined knowledge-editing method designed to improve editing efficacy while better preserving pre-trained capabilities. This approach differs from raw PSFT that updates the whole LLMs, we restrict updates exclusively to specific FFN layers, widely recognized as the primary stores of factual knowledge (Meng et al., 2022a). This targeted update strategy effectively injects new knowledge without modifying the model architecture, while minimizing disruption to pre-trained capabilities.
We consider a knowledge editing instance , where is the input query and is the target answer. Let parameterized by denote the vanilla LLM before the current edit. We partition the current parameters into two disjoint sets: the trainable FFN parameters and the frozen parameters , such that . The objective of TPSFT is to optimize to learn the updated parameters , yielding the post-edit policy where .
A distinctive aspect of TPSFT is the use of self-generated CoT-augmented training labels. For each instance , we: (1) prompt the vanilla LLM to generate a CoT reasoning path for (see Appendix A.5), and (2) replace the generated answer with , yielding the training target . This design integrates the CoT into the supervision signal. Rather than overfitting to a sharp, isolated input-output mapping, the model learns to associate the new knowledge with valid reasoning patterns. To ensure data quality, we filter out generated chains that are logically incompatible with the target (see Appendix A.5), thereby preserving the model’s intrinsic generation capabilities while injecting the new fact.
Formally, given the editing dataset , for each token position in the target sequence , we define the state and the target token . We optimize by minimizing the following TPSFT loss:
@@
Here, is a hyperparameter that defines the clipping threshold. Following Zhu et al. (2025), the ratio clipping mechanism provides implicit trust-region regularization by (i) bounding the log-probability change per sample to , and (ii) nullifying gradients when the ratio exceeds the upper bound. The probability ratio is the core of this objective and is defined as:
@@
where is the probability of generating the target token given the context from the model with the current FFN parameters, and is the corresponding probability from the reference policy. At the start of the editing process, this reference policy is the initial vanilla LLM. For each subsequent edit instance in the sequential editing process, it is then updated to be the state of the model resulting from the immediately preceding edit.
This objective function acts as a trust-region regularizer via ratio clipping. The term encourages the model to increase the likelihood of the target answer, similar to standard supervised fine-tuning. Crucially, when this ratio exceeds the threshold , the gradient with respect to is nullified. This design effectively removes optimization pressure from well-learned samples, constraining parametric deviation and preventing individual edits from dominating the parameter updates.
However, despite these constraints, TPSFT remains a teacher-forced training mechanism implicitly conditioned on ground-truth prefixes . During inference, the model must rely on its own generated history . This discrepancy creates a fundamental train-inference mismatch: while the knowledge is parametrically encoded, it may not be reliably elicited during autoregressive generation. This critical limitation necessitates the subsequent consolidation phase.
4.2 Knowledge Consolidation via GRPO
As motivated in § 3.2, parametric editing alone is insufficient—edited knowledge must be consolidated at the behavioral level. Building on GRPO’s trajectory-level optimization (§ 3.1), we design a consolidation algorithm that transforms parametrically encoded knowledge into reliably generated outputs.
Starting from the edited model obtained via TPSFT, we optimize the policy (initialized as ) through autoregressive sampling on the editing dataset . For each query , the model generates complete trajectories , which are then scored against the target answer . The group-relative advantage formulation enables the model to learn from comparative trajectory quality without requiring absolute reward calibration. The reference policy is set to , ensuring that consolidation refines generation behavior while preserving the parametric edits from Stage I.
| ZsRE | COUNTERFACT | QAEdit | |||||||||||
| Method | Reli. | Gen. | Loc. | Avg. | Reli. | Gen. | Loc. | Avg. | Reli. | Gen. | Loc. | Avg. | |
| Qwen2.5-7B -Instruct | Pre-edit | 4.4 | 3.2 | 28.5 | 12.0 | 1.0 | 0.5 | 36.9 | 12.8 | 9.8 | 10.1 | 36.2 | 18.7 |
| FT-M | 5.6 | 5.5 | 23.1 | 11.4 | 3.2 | 3.1 | 24.4 | 10.2 | 14.6 | 14.5 | 30.7 | 19.9 | |
| MEMIT | 0.0 | 0.1 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.1 | 0.4 | 0.3 | 0.2 | 0.3 | |
| ALPHAEDIT | 15.9 | 11.5 | 6.8 | 11.4 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| WISE | 4.5 | 3.3 | 19.1 | 9.0 | 1.4 | 1.5 | 31.0 | 11.3 | 7.1 | 9.7 | 16.9 | 11.2 | |
| GRACE | 4.8 | 4.7 | 28.1 | 12.5 | 1.9 | 1.2 | 36.8 | 13.3 | 14.2 | 13.0 | 36.0 | 21.1 | |
[…]
| Llama-3 -8B-Instruct | Pre-edit | 2.8 | 2.4 | 38.6 | 14.6 | 0.6 | 0.8 | 31.8 | 11.1 | 12.7 | 12.5 | 44.3 | 23.2 |
| FT-M | 16.6 | 15.5 | 29.3 | 20.5 | 27.9 | 18.6 | 10.5 | 19.0 | 34.1 | 33.2 | 30.1 | 32.5 | |
| MEMIT | 0.1 | 0.1 | 0.0 | 0.1 | 0.3 | 0.7 | 0.4 | 0.5 | 0.2 | 0.7 | 0.0 | 0.3 | |
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | 13.0 | 61.0 | 43.8 | 16.1 | 40.3 | 18.2 | 14.9 | 7.5 | 13.5 | |
| WISE | 4.3 | 3.1 | 2.2 | 3.2 | 1.3 | 0.8 | 31.3 | 11.1 | 8.1 | 13.3 | 0.9 | 7.4 | |
| GRACE | 5.5 | 3.9 | 37.9 | 15.8 | 2.1 | 0.7 | 32.1 | 11.6 | 15.4 | 13.4 | 34.2 | 21.0 | |
[…]